1. Consider the family of differential equations dy/dx = y^3 + ky + k^3 .

Please Help me with it, thanks so much
Answer #1

3 dr yot kytk² Solution - We have given problem, dy a) we have Identify are there any equilbrium 8012 where k=0? If so, WhatC) we have to draw the phare diagram for when R = 1/2 dy dx y²tky tk² =) Y ²- ty - to y (Y²_ ² ) - te =0, ² t = 0 y2 – y =Y²x - byx~tx=7 y = y3 x 1 x - 13 tx Now lim y300 - Lyoo limxta to Hence lim xao does not exsit. For which value of k do these